# A Survey of Language-Based Communication in Robotics

### v1

@@ 3.1.1. Task Breakdown and Planning
The first and most straightforward strategy is for the human to define a high-level mission objective which is broken down by the model into subtasks. This approach can generate plans which directly inform robot actions, although typically via a proxy such as labelled action policies. The most famous example is SayCan, which uses an LLM to suggest a suitable action in response to a textual prompt, and then selects the best actions from a list of pretrained affordances (Brohan et al., 2023b). Other works use various models to reduce the environment and interaction to a textual space; perceptions are described in text by naming objects and features in the environment. An LLM then selects which object to choose, and other models perform robotic movement and action (Dorbala et al., 2023; Wu et al., 2023d). This can be used to respond to direct human instructions, or to anticipate human needs from implicit queries (Kakodkar et al., 2023). Oftentimes instructions from humans are incidental and unstructured and it is important to be able to understand and act upon these (Tang et al., 2023), or else an ambiguity exists which can only be resolved by considering the current situation alongside the human’s instruction (Zhi-Xuan et al., 2024). Plans such as this can also be generated from a combination of task description and current state of the robot (Ha et al., 2023). It is possible to “freeze” the LLM; place it into the architecture and train other models around it without altering the LLM’s parameters. Smaller models can then be used to adapt what the frozen LLM outputs as a plan such that it can better be understood and performed by the robot (Huang et al., 2022a). An advantage of this approach is that it scales well. The process should theoretically be the same whether the lowest level tasks are atomic or are still quite high level. Large environments can be operated in with large sets of tasks handled by a system of this type, with the planning agent solving numerous subtasks iteratively towards the initial goal (Rana et al., 2023). A single LLM can also be used to model a team of robots and divide work between them based on human input. For example, a high level task can be broken down and assigned to robots based on their available skills (Kannan et al., 2023).
@@
An interesting variant of this approach takes a textual input and uses it to generate an imagined image of what the end goal should look like. Employing the DALL-E (Ramesh et al., 2021) image generation model, this method is following a similar process, but by using an image as its immediate goal, it is potentially describing the task in a manner which is closer to its perception. The concrete problem of properly understanding the task in text form is here moved slightly to a new form (as of course the generated image by the model may not align with the human’s intention). A useful advantage of this approach however is that the goal as the robot understands it is presented to the user at the start of execution, not the end. This may give the human more time to intervene in the likely event that the machine has misunderstood (Kapelyukh et al., 2023). Another means of utilising the LLM outside of the core control loop of the robot is to generate simulation data on which the robot can train. This is similar in that it fundamentally generates the conditions for the robot, and another model such Reinforcement Learning (RL), with the likely advantage of more explainability and predictability, can be used to actually perform he execution (Wang et al., 2023a). The same idea can also be used to generate alternative prompts and images for a more end-to-end approach (Yu et al., 2023b). Trajectories can also be learned and imagined from textual inputs and initial environment state, which are then used to define the inputs to a standard robotics transformer model (Gu et al., 2023a)
3.1.2. Writing Code and Defining Rewards
@@
Although a key advantage of using language is to replace the requirement for a user to understand and engage with technical command and code, a viable approach is to use the LLM as a translator. A human can define a behaviour in natural language, and the agent uses this to write code for the robot. Such code can be written to define a reward function for an RL agent. This utilises the twin advantages of LLMs for interpreting human commands, and RL for training the robot to achieve the behaviour (Yu et al., 2023a). This approach allows the LLM to both train the model and to incrementally update the reward function based on current performance (Ma et al., 2023) and, if available, visual feedback (Narin, [n. d.]). It can also help to build a policy that functions well, but can be steered by human input towards one of several acceptable options (Hu and Sadigh, 2023). A similar approach is to write evaluation methods for labelling successful or unsuccessful trials (Ha et al., 2023). Alternatively, an LLM can directly write code to manipulate the robot, just as a human would. By giving the model an API of functions with descriptions, it can write a new method to control the robot around comments using a library of provided functions (Singh et al., 2023; Huang et al., 2023a; Liang et al., 2023), or can directly break down an objective and write the method with guidance on syntax (Hu et al., 2024). This kind of approach is sometimes augmented with other models to empower it, such as using Monte-Carlo Tree Search to find the best solution which is partially constructed and judged by a transformer model (Zhang et al., 2023b). A key point of distinction between these approaches is how the environment and action affordances are presented to the robot. A popular solution is to provide example code in libraries, but fake methods can also be used, for example the robot is given a method to detect objects which outputs a set of interactable elements in front of it; by presenting the information to the system in this manner it naturally flows into the coding environment and allows the robot to consider it in the same way it would any variable from its corpus of coding examples. High level abstractions like this also allow pretrained or manually written skills to be treated as a callable method without requirement for the robot to engage with the specifics of how this is performed (Lin et al., 2023a). The input space in such examples can be produced with segmented maps of objects in the environment, and other inputs such as gestures (Lin et al., 2023b).
3.1.3. Direct Integration of Human Input
@@
LLMs or language embeddings can be inserted directly into the model in a manner where a human could not feasibly replace the LLM. Instead of using the language in a logical, sequential way as with the other categories, this approach is more “black-box” like; placing some text or a textual embedding into part of the model or its input, and training other models around this. Fundamentally, this still means that the language is used to instruct and command the robot, but not in an explainable or interpretable way. Some works place the LLM directly into the model for the robot, such as by using a combination of image labelling and extracted keywords from a human command to recognise landmarks during navigation (Shah et al., 2023). This sometimes demands a more careful model which considers that human input is sometimes incorrect, and this should be recognised and adapted to wherever possible (Taioli et al., 2024). LLMs frequently form part of the action selection process, such as mapping human instructions to the most similar action generated from environmental observations (Yow et al., 2024). Other approaches use a single LLM to process the instruction alongside representations of the environment perception or robot state (Ni et al., 2023; Zhu et al., 2024a). Human input can also be fed into an LLM to adapt the inputs to another model (Wen et al., 2024), even another, frozen, LLM (Bärmann et al., 2023). Approaches of this form facilitate an updating control loop if required, where the model is updated and trained based on success, failure, or human alterations and corrections (Liu et al., 2023c). This type of structure could be promising in a lifelong learning context where robots can learn from their mistakes and correct undesired behaviours or shortcomings in the future. The LLM itself can even be called upon to generate better prompts for its own model to train on after failure (Bing et al., 2023). Already functional behaviours can be also be subtly modified with human alterations by appending a representation of a command relative to an object of interest, such as instruction the robot to “keep a much bigger distance from the computer” when delivering drinks to a human (Bucker et al., 2023). Similarly, a policy can be trained that alters actions given by a human on a handheld controller, combining this control with a textual instruction and outputting joint parameters that satisfy both (Karamcheti et al., 2022). These approaches allow a human to subtly improve the performance of an already trained and functional model without requirement for further training or alteration (Shi et al., 2024). This method of altering robot behaviour may be highly practical for non-technical users to work alongside robots as it may not be practical to finetune a robot to a specific task or environment; instead having the users advise it to subtly change using language could be much more feasible in future industry (Cui et al., 2023b; Huang et al., 2023b). Another approach is to use behavioural cloning to learn from humans tele-operating the robot while narrating their tasks. This allows the policy to factor a linguistic instruction into the input space of the model during training, so that the policy can be used to control the robot with spoken commands during execution (Lynch et al., 2023)
3.2. Robot to Human Communication
@@
Agents are often difficult to understand for human operators and bystanders (Rosenfeld and Richardson, 2019), and a sensible application of LLMs is to address this problem. Robots can use language to describe their actions, beliefs, and intentions to a human; such an approach removes the requirement for expert users who can debug the robot’s beliefs, instead communicating in a way that almost all can understand — language. The generative aspect of LLMs is naturally suited to this task, as well as the question-answering mode that is often employed (Wu et al., 2017). This is especially relevant when paired with visual models and/or in safety-critical scenarios; describing why an agent took a certain action could help operators catch mistakes when human lives are at risk. The “Robot to Human Communication” category describes work which employs LLMs to feed information back to humans for purposes ranging from explainability to concisely representing their observations of the environment.
@@
Table 2. A selection of works using language for Human-to-Robot tasking. The “LLM” column indicates the most recent LLM they used (not all if they used several variants). “Language used to” is left high level to allow papers to be further subgrouped.
Modality Paper LLM Language used to
Explainability Wayve LINGO-1 Describe AV actions & Q-A
Zhang et al. GPT-4 Describe agent actions & Q-A
González-Santamarta et al. GPT-3.5 & Alpaca Describe behaviour through log files
Chen et al. GPT-3.5 Describe AV actions
Liu et al. GPT-4 Describe task from video
Lu et al. GPT-3.5 Q-A in terms of reward
Raman et al. Syntax trees Describe feasibility
Gong and Zhang N/A (templates) Signal intentions
Garcia et al. N/A (templates) Q-A
Das et al. Syntax trees Diagnose errors and suggest resolution
Asking for Help Ren et al. PaLM-2L & GPT-3.5 Ask when certainty is low
Mullen Jr and Manocha GPT-4 Ask when certainty is low
Hori et al. GPT-3.5 Ask to resolve ambiguity
Padmakumar et al. Episodic Transformer Ask to resolve ambiguity
Doğan et al. N/A Ask to resolve ambiguity
Banerjee et al. LSTM Describe env. & ask for guidance
Sattar and Dudek N/A Ask when certainty is low
Song et al. GPT-3 Ask to resolve ambiguity
Thomason et al. N/A Ask to resolve ambiguity
Zhang et al. GPT-3.5 Ask to guide through decision tree
Wang et al. GPT-3.5 Ask to resolve ambiguity
Zheng et al. BART Ask to resolve ambiguity
3.2.1. Explainability
@@
Typically, language models which are employed for robot-human communication will simply output messages one-way to the user. This may take the form of the agent logging its actions and decisions, and using language models to efficiently communicate the relevant parts of this data. Alternatively, question answering models can be connected to the agent to provide the user with a direct interface to the system. The user can now query the system about its reasoning for certain actions or which parts of the environment it is currently paying most attention to.
@@
Early works (before LLMs were commonplace) used predefined structures such as syntax trees to define text based outputs to the human (Raman et al., 2013). This may take the form of hand-written labels based on common outputs such as failure messages (Das et al., 2021), or templates which are adapted to generate explanations of behaviour (Gong and Zhang, 2018). Answers to common debugging questions could also be generated to aid the human in understanding (Garcia et al., 2018). The overall theme from this area, however, was that there was a need for better verbal, language-based cues and explanations (Han et al., 2021).
@@
Since the introduction of LLMs, much more work has appeared in this space. LLMs can be used to analyse robot log files and explain their behavior (González-Santamarta et al., 2023; Sobrín-Hidalgo et al., 2024), or they can answer questions in terms of the agent’s reward function (Lu et al., 2023). Alternatively, complex descriptions of the reasoning behind actions can be generated in a multi-agent setting (Zhang et al., 2023d). This can include multiple sources of data such as by captioning audiovisual information to describe the solving of a task (Liu et al., 2023a). A promising application of this concept is explaining the behaviour and decision making of autonomous vehicles (Chen et al., 2023d), either by narrating as the vehicle drives, or by using a question-answer prompt system to allow a conversation with the agent about what it is focussing on and how it intends to act (Wayve, 2023). The dialogical approach, whereby the use can actively ask the model questions, may be a more natural form of explainability that extends from a language-based instruction method (Nwankwo and Rueckert, 2024)
3.2.2. Asking for Human Assistance
@@
Into the robot-human category we also place robots that use the human to support them on request. These systems generally plan their actions and output them to a human who can interrupt if a robot made a mistake. The key driver of this approach is deciding when to ask for the human to help; it is imprudent to flood a human operator with questions and requests because they become overloaded. Approaches of this type typically compute some kind of confidence score and use that to determine whether they should proceed (Sattar and Dudek, 2011; Mullen Jr and Manocha, 2024). The system could also describe the problem it is facing and ask for human instruction (Banerjee et al., 2021; Thomason et al., 2020). A more sophisticated approach is to determine the ambiguity which is causing the robot to be unsure. This is a natural behaviour in humans to clarify by asking a follow-up question. This can take the form of asking for additional details about the human’s command, and the details of what they want to be achieved (Hori et al., 2024; Zhang et al., 2023e; Zheng et al., 2022), or could be used to resolve simple issues that the robot is struggling with such as location or task status (Padmakumar et al., 2022). Robots which resolve ambiguities by asking a question which rephrases the task or specifically targets the area of difficulty (for example asking “is the object you want to the left of X?” outperform those which ask for a simple rewording of the same task (Doğan et al., 2022). A well-designed approach of this type will highlight to the user the issue, or will provide a possible point of reference such as “I cannot find X, but I can see Y” so that the user can confirm how it should proceed (Song et al., 2023). This has implications on safety; robots can ask a human for assistance if they are not sure what the safest acceptable course of action might be (Wang et al., 2024b). Finally, robots can frame their question as more targeted than a yes or no answer. In a scenario where ambiguity is present, the robot can choose those optional actions with highest confidence and ask for clarification between them directly to the user (Ren et al., 2023). This allows the human to more quickly resolve the exact issue with a specific command instead of giving a yes or no answer that might devolve into further issues.
3.3. Robot to Robot Communication

### v3

@@ 3.1.1 Task Breakdown and Planning
The first and most straightforward strategy is for the human to define a high-level mission objective which is broken down by the model into subtasks. This approach can generate plans which directly inform robot actions, although typically via a proxy such as labelled action policies. The most famous example is SayCan, which uses an LLM to suggest a suitable action in response to a textual prompt, and then selects the best actions from a list of pretrained affordances Brohan et al., 2023b . Other works use various models to reduce the environment and interaction to a textual space; perceptions are described in text by naming objects and features in the environment. An LLM then selects which object to choose, and other models perform robotic movement and action Dorbala et al., (2023); Wu et al., 2023d . This can be used to respond to direct human instructions, or to anticipate human needs from implicit queries Kakodkar et al., (2023). Oftentimes instructions from humans are incidental and unstructured and it is important to be able to understand and act upon these Tang et al., (2023), or else an ambiguity exists which can only be resolved by considering the current situation alongside the human’s instruction Zhi-Xuan et al., (2024). Plans such as this can also be generated from a combination of the task description and the current state of the robot Ha et al., (2023). It is possible to “freeze” the LLM; place it into the architecture and train other models around it without altering the LLM’s parameters. Smaller models can then be used to adapt what the frozen LLM outputs as a plan such that it can better be understood and performed by the robot Huang et al., 2022a . An advantage of this approach is that it scales well. The process should theoretically be the same whether the lowest level tasks are atomic or are still quite high-level. Large environments can be operated in with large sets of tasks handled by a system of this type, with the planning agent solving numerous subtasks iteratively towards the initial goal Rana et al., (2023). A single LLM can also be used to model a team of robots and divide work between them based on human input. For example, a high-level task can be broken down and assigned to robots based on their available skills Kannan et al., (2023).
@@
An interesting variant of this approach takes textual input and uses it to generate an imagined image of what the end goal should look like. Employing the DALL-E Ramesh et al., (2021) image generation model, this method follows a similar process, but by using an image as its immediate goal, it is potentially describing the task in a way closer to its perception. The concrete problem of properly understanding the task in text form is here moved slightly to a new form (as of course the generated image by the model may not align with the human’s intention). A useful advantage of this approach however is that the robot’s understanding of the goal is presented to the user at the start of execution, not the end. This may give the human more time to intervene in the likely event that the machine has misunderstood or hallucinated Kapelyukh et al., (2023). Another means of utilising the LLM outside of the core control loop of the robot is to generate simulation data on which the robot can train. This is similar in that it fundamentally generates the conditions for the robot, and another model such Reinforcement Learning (RL), with the likely advantage of more explainability and predictability, can be used to actually perform the execution Wang et al., 2023d . The same idea can also be used to generate alternative prompts and images for a more end-to-end approach Yu et al., 2023a . Trajectories can also be learned and imagined from textual inputs and initial environment state, which are then used to define the inputs to a standard robotics transformer model Gu et al., 2023a .
3.1.2 Writing Code and Defining Rewards
@@
Although a key advantage of using language is to replace the requirement for a user to understand and engage with technical command and code, a viable approach is to use the LLM as a translator. A human can define a behaviour in natural language, and the agent uses this to write code for the robot. Such code can be written to define a reward function for an RL agent. This utilises the twin advantages of LLMs for interpreting human commands, and RL for training the robot to achieve the behaviour Yu et al., 2023b . This approach allows the LLM to both train the model and to incrementally update the reward function based on current performance Ma et al., (2023) and, if available, visual feedback Narin, (2024). It can also help to build a policy that functions well, but can be steered by human input towards one of several acceptable options Hu and Sadigh, (2023). A similar approach is to write evaluation methods to label successful or unsuccessful trials Ha et al., (2023). Alternatively, an LLM can directly write code to manipulate the robot, just as a human would. By giving the model an API of functions with descriptions, it can write a new method to control the robot based on human-written comments and using a library of provided functions Singh et al., (2023); Huang et al., 2023a ; Liang et al., (2023), or can directly break down an objective and write the method with syntax guidance Hu et al., (2024). This kind of approach is sometimes augmented with other models to empower it, such as using Monte-Carlo Tree Search (MCTS) to find the best solution which is partially constructed and judged by a transformer model Zhang et al., 2023e . A key point of distinction between these approaches is how the environment and action affordances are presented to the robot. A popular solution is to provide example code in libraries, but fake methods can also be used, for example, the robot is given a method to detect objects which outputs a set of interactable elements in front of it; by presenting the information to the system in this manner, it naturally flows into the coding environment and allows the robot to consider it in the same way it would any variable from its corpus of coding examples. High-level abstractions like this also allow pretrained or manually written skills to be treated as a callable method without a requirement for the robot to engage with the specifics of how this is performed Lin et al., 2023a . The input space in such examples can be produced with segmented maps of objects in the environment, and other inputs such as gestures Lin et al., 2023b .
3.1.3 Direct Integration of Human Input
@@
LLMs or language embeddings can be inserted directly into the model in a manner where a human could not feasibly replace the LLM. Instead of using the language in a logical, sequential way as with the other categories, this approach is more “black-box” like; placing some text or a textual embedding into part of the model or its input, and training other models around this. Fundamentally, this still means that the language is used to instruct and command the robot, but not in an explainable or interpretable way. Some works place the LLM directly into the model for the robot, such as by using a combination of image labelling and extracted keywords from a human command to recognise landmarks during navigation Shah et al., (2023). This sometimes demands a more careful model which considers that human input is sometimes incorrect, and this should be recognised and adapted to wherever possible Taioli et al., (2024). LLMs frequently form part of the action selection process, such as mapping human instructions to the most similar action generated from environmental observations Yow et al., (2024). Other approaches use a single LLM to process the instruction alongside representations of the environment perception or robot state Ni et al., (2023); Zhu et al., 2024a . Human input can also be fed into an LLM to adapt the inputs to another model Wen et al., (2024), even another, frozen, LLM Bärmann et al., (2023). Approaches of this form facilitate an updating control loop if required, where the model is updated and trained based on success, failure, or human alterations and corrections Liu et al., 2023a . This type of structure could be promising in a lifelong learning context where robots can learn from their mistakes and correct undesired behaviours or shortcomings in the future. The LLM itself can even be called upon to generate better prompts for its own model to train on after failure Bing et al., (2023). Already functional behaviours can also be subtly modified with human alterations by appending a representation of a command relative to an object of interest, such as instructing the robot to “keep a much bigger distance from the computer” when delivering drinks to a human Bucker et al., (2023). Similarly, a policy can be trained that alters actions given by a human on a handheld controller, combining this control with a textual instruction and outputting joint parameters that satisfy both Karamcheti et al., (2022). These approaches allow a human to subtly improve the performance of an already trained and functional model without a requirement for further training or alteration Shi et al., (2024). This method of altering robot behaviour may be highly practical for non-technical users to work alongside robots as it may not be practical to finetune a robot to a specific task or environment; instead having the users advise it to subtly change using language could be much more feasible in future industry Cui et al., 2023b ; Huang et al., 2023b . Another approach is to use behavioural cloning to learn from humans tele-operating the robot while narrating their tasks. This allows the policy to factor a linguistic instruction into the input space of the model during training, so that the policy can be used to control the robot with spoken commands during execution Lynch et al., (2023).
3.2 Robot to Human Communication
@@
Agents are often difficult to understand for human operators and bystanders Rosenfeld and Richardson, (2019), and a sensible application of LLMs is to address this problem. Robots can use language to describe their actions, beliefs, and intentions to a human; such an approach removes the requirement for expert users who can debug the robot’s beliefs, instead communicating in a way that almost all can understand — language. The generative aspect of LLMs is naturally suited to this task, as well as the question-answering mode that is often employed Wu et al., (2017). This is especially relevant when paired with visual models and/or in safety-critical scenarios; describing why an agent took a certain action could help operators catch mistakes when human lives are at risk. The “Robot to Human Communication” category describes work which employs LLMs to feed information back to humans for purposes ranging from explainability to concisely representing their observations of the environment.
@@
Table 2: A selection of works using language for Robot-to-Human feedback. The “LLM” column indicates the most recent LLM they used (not all if they used several variants). “Application” is left high level to allow papers to be further subgrouped.
Modality Paper LLM Application Explainability Wayve, (2023) LINGO-1 Describing actions Zhang et al., 2023f GPT-4 Describing actions & Question-Answering González-Santamarta et al., (2023) GPT-3.5 & Alpaca Describing behaviour Chen et al., 2023b GPT-3.5 Describing actions Liu et al., 2023c GPT-4 Describing task Lu et al., (2023) GPT-3.5 Question-Answering Raman et al., (2013) Syntax trees Describing feasibility Gong and Zhang, (2018) N/A (templates) Signaling intentions Garcia et al., (2018) N/A (templates) Question-Answering Das et al., (2021) Syntax trees Diagnosing errors Asking for Help Ren et al., (2023) PaLM-2L & GPT-3.5 Asking when certainty is low Mullen Jr and Manocha, (2024) GPT-4 Asking when certainty is low Hori et al., (2024) GPT-3.5 Asking to resolve ambiguity Padmakumar et al., (2022) Episodic Transformer Asking to resolve ambiguity Doğan et al., (2022) N/A Asking to resolve ambiguity Banerjee et al., (2021) LSTM Describing env. & asking for guidance Sattar and Dudek, (2011) N/A Asking when certainty is low Song et al., (2023) GPT-3 Asking to resolve ambiguity Thomason et al., (2020) N/A Asking to resolve ambiguity Zhang et al., 2023g GPT-3.5 Asking to guide through decision tree Wang et al., 2024a GPT-3.5 Asking to resolve ambiguity Zheng et al., (2022) BART Asking to resolve ambiguity Liang et al., 2024a GPT-4 Asking to resolve ambiguity
3.2.1 Explainability
@@
Typically, language models which are employed for robot-human communication will simply output messages one-way to the user. This may take the form of the agent logging its actions and decisions, and using language models to efficiently communicate the relevant parts of this data. Alternatively, question-answering models can be connected to the agent to provide the user with a direct interface to the system. The user can now query the system about its reasoning for certain actions or which parts of the environment it is currently paying the most attention to.
@@
Early works (before LLMs were commonplace) used predefined structures such as syntax trees to define text-based outputs to the human Raman et al., (2013). This may take the form of hand-written labels based on common outputs such as failure messages Das et al., (2021), or templates which are adapted to generate explanations of behaviour Gong and Zhang, (2018). Answers to common debugging questions could also be generated to aid the human in understanding Garcia et al., (2018). The overall theme from this area, however, was that there was a need for better verbal, language-based cues and explanations Han et al., (2021).
@@
Since the introduction of LLMs, much more work has appeared in this space. LLMs can be used to analyse robot log files and explain their behavior González-Santamarta et al., (2023); Sobrín-Hidalgo et al., (2024), or they can answer questions in terms of the agent’s reward function Lu et al., (2023). Alternatively, complex descriptions of the reasoning behind actions can be generated in a multi-agent setting Zhang et al., 2023f . This can include multiple sources of data such as by captioning audiovisual information to describe the solving of a task Liu et al., 2023c . A promising application of this concept is explaining the behaviour and decision-making of autonomous vehicles Chen et al., 2023b , either by narrating as the vehicle drives, or by using a question-answer prompt system to allow a conversation with the agent about what it is focussing on and how it intends to act Wayve, (2023). The dialogical approach, whereby the user can actively ask the model questions, may be a more natural form of explainability that extends from a language-based instruction method Nwankwo and Rueckert, (2024).
3.2.2 Asking for Human Assistance
@@
Into the robot-human category, we also place robots that use the human to support them on request. These systems generally plan their actions and output them to a human who can interrupt if a robot made a mistake. The key driver of this approach is deciding when to ask for the human to help; it is imprudent to flood a human operator with questions and requests because they become overloaded. Approaches of this type typically compute some kind of confidence score and use that to determine whether they should proceed Sattar and Dudek, (2011); Mullen Jr and Manocha, (2024). The system could also describe the problem it is facing and ask for human instruction Banerjee et al., (2021); Thomason et al., (2020). A more sophisticated approach is to determine the ambiguity which is causing the robot to be unsure. This is a natural behaviour in humans to clarify by asking a follow-up question. This can take the form of asking for additional details about the human’s command, and the details of what they want to be achieved Hori et al., (2024); Zhang et al., 2023g ; Zheng et al., (2022), or could be used to resolve simple issues that the robot is struggling with such as location or task status Padmakumar et al., (2022). Robots which resolve ambiguities by asking a question which rephrases the task or specifically targets the area of difficulty (for example asking “is the object you want to the left of X?” outperform those which ask for a simple rewording of the same task Doğan et al., (2022). A well-designed approach of this type will highlight the issue to the user, or will provide a possible point of reference such as “I cannot find X, but I can see Y” so that the user can confirm how it should proceed Song et al., (2023). This has implications on safety; robots can ask a human for assistance if they are not sure what the safest acceptable course of action might be Wang et al., 2024a . Finally, robots can frame their question as more targeted than a yes or no answer. In a scenario where ambiguity is present, the robot can choose those optional actions with the highest confidence and ask for clarification between them directly to the user Ren et al., (2023), or else present some candidate actions from which the human selects Liang et al., 2024a . This allows the human to more quickly resolve the exact issue with a specific command instead of giving a yes or no answer that might devolve into further issues.
3.3 Robot to Robot Communication